# Foundational GPT Model for MEG

### v1

@@ 2.3 HMM statistics of generated data match real data
Next, we assessed how well the fitted models can recursively generate data with the same spatial, temporal and spectral multi-channel characteristics as real MEG data. HMMs are an established way for doing unsupervised discovery of multi-channel dynamics in real neuroimaging data, and have been used to characterise the spatial, temporal and spectral characteristics of brain networks in MEG data (Rabiner,, 1989; Vidaurre et al., 2018a, ).
Separately, we trained each model on a single sample subject (with condition embeddings), generated multi-channel data, and then inferred a 12-state HMM, with each state modelling the multi-channel data as a multi-variate Normal (MVN) distribution. The number of states was chosen based on previous work (Vidaurre et al., 2018a, ). Example state time-courses generated from all models are plotted in Figure 2 to qualitatively illustrate the differences in the generated dynamics. Note that since a separate HMM is trained each time, the states are not matched between models or with the real data. States could be matched post-HMM training by looking for similarity in the observation models, but here we opted to plot unmatched state time-courses. Since each generated timeseries is different, comparison across state timecourses would not be useful anyway.
Figure 2: Example state timecourses from the HMMs trained on each model’s generated data (rows). Each state is represented by a different colour. Note that state indices and timecourses are not matched across models.
@@
Figure 3: Distributions across the 12 states inferred by an HMM using multi-channel data generated from the different models, or using the real sensor space MEG data (columns). Four different summary statistics are shown describing the state dynamics (rows).
@@
We extracted four summary statistics on the inferred state timecourses and compared their distribution over states. These statistics are shown across models in Figure 3, alongside those for an HMM trained on the real multi-channel MEG data. Across the four summary statistics we can see that the real data has high variance in the distribution over states. AR(255) and WavenetFullChannelMix fail to produce data with variable state statistics, and even the mean over states is not captured well. WavenetFullChannel does a great job at capturing the mean of the state distributions, but still produces data with relatively invariant states. ChannelGPT2 seems to best capture the distributions across all four statistics, especially for the mean interval and switching rate. This shows that Transformer-based models can generate data that better matches the HMM-inferred dynamics of real MEG data.
In addition to state statistics, we also computed the power spectra of each state across the timeseries. In MEG data different states might capture oscillatory activity with specific frequencies. The extracted power spectra from the inferred state time courses is shown in Figure 4. We can see that the HMM trained on the MEG data contains many states that capture the 10 Hz peak, with fewer states having a 20 Hz peak. It is also clear that the states of the HMM fitted to the WavenetFullChannelMix generated timeseries do not contain these spectral peaks. While the AR(255) does contain states with a 10 Hz peak, the shape does not match the data well, and also states do not show the same variability as in real data.
In contrast, ChannelGPT2, matches the state PSDs of the real data very well, further demonstrating the superiority of Transformer models in capturing complex neural dynamics. While WavenetFullChannel also improves substantially over the AR(255) power spectra, it falls short in capturing the 20 Hz peak and the heterogeneity between states observed in the real data and the generated data of ChannelGPT2. This and previous analyses show that the combination of channel-independence and a Transformer-based architecture are critical for matching the dynamics of real data.
(a) Data
(b) AR(255)
(c) WFCM
(d) WavenetFullChannel
(e) ChannelGPT2
Figure 4: Power spectral density of HMM states inferred on the generated data of each model and on real MEG data. WFCM refers to WavenetFullChannelMix. Each line is the PSD of a different state. Note that states are not matched across models. Horizontal axis represents frequency in Hz.
2.4 Evoked activity in generated data matches real data
@@
Figure 5: Comparison of evoked timecourses of 2 channels across our task-conditioned models. Stimulus onset is at 0 seconds and offset is at 500ms. The peak occurring after 500 ms indicates a visual response to the offset of the stimulus (removal of the image). Shading indicates variability across trials.
The analyses in the previous section considered metrics for assessing the ability of the candidate foundation models to generate timeseries without requiring any a priori knowledge of the timing of brain activity. Here, we use knowledge of the experimental task timings in the Cichy et al., (2016) data, to provide insight into the ability of the foundational models to generate realistic task data.
@@
By simple epoching of the generated timeseries based on the known task labels, we can compute evoked responses in the data generated by our models. We do this for all models except AR(255) as it did not include task labels in its model. To compare the shape of average evoked responses, we averaged over all epochs in both real data and the generated timeseries. This results in data of shape $\bar{\mathbf{X}}\in\mathbb{R}^{C\times T}$ where $C=306$ is the number of channels and $T=1000$ ms is the trial/epoch length.
@@
The evoked responses across our models and the real data in a frontal and a visual channel are shown in Figure 5. While Wavenet models completely fail to capture the evoked time-course, ChannelGPT2 does a remarkably good job, especially in the visual channel. This is not surprising as the dataset is collected from a visual experiment, so most activity is visual. ChannelGPT2 closely matches both the amplitude and the timing of the evoked response peaks across the whole 1-second epoch. Variability across trials is also well matched.
(a) WFC
(b) WFCM
(c) ChannelGPT2
Figure 6: Correlation between the time-courses of the mean (over individual epochs) evoked responses from the real data and mean evoked responses generated by each model. The correlation values are visualised across sensors. WFC refers to WavenetFullChannel and WFCM refers to WavenetFullChannelMix. Darker reds indicate higher correlation.
(a) WFC
(b) WFCM
(c) ChannelGPT2
Figure 7: Correlation between the time-courses of the variance (over individual epochs) of the mean evoked responses from the real data and the variance of the mean evoked responses generated by each model. The correlation values are visualised across sensors. WFC refers to WavenetFullChannel and WFCM refers to WavenetFullChannelMix. Darker reds indicate higher correlation.
Figure 8: Evoked response state timecourses of HMMs trained on the real MEG data and generated data from each of our task-conditioned models. Note that states are not matched between models. Image presentation starts at 0 seconds and ends at 0.5 seconds.
As expected, ChannelGPT2 generates data with evoked responses that have much higher correlation with evoked responses from real data, and slightly higher correlation in visual areas compared to other channels, matching the known topography of visual evoked responses. In other models the correlation is low, and spatially better in frontal areas, likely because the evoked responses here are noisier providing an easier fit.
Figure 7 shows the correlation between the variance (over individual epochs) time-courses of the mean evoked response obtained from the actual data and the evoked responses obtained from data generated by each model. This captures a measure of the ability of the models to represent the trial-to-trial variability found in the real data. Again, ChannelGPT2 generates data that has the highest correlations with the real data, with higher values in channels in the back of the head, appropriately capturing the topography of response variability. Other models have similar spatial distribution, and notably WavenetFullChannel also produces evoked responses with variance partially matching the real data.
Finally, a different way to assess task-related activity is to examine the evoked state time-courses from the HMMs fitted on the real and model generated timeseries. Rather than looking at individual channels, this provides an overall view of which HMM state gets activated when, during individual trials. This is computed by simply epoching the state timecourse, and averaging over all trials. We plot these for the real data and each generated timeseries in Figure 8. As expected, the HMM trained on models other than ChannelGPT2 shows poor evoked state time-courses. ChannelGPT2 generated data produces states with similar evoked dynamics and variability as the real data.
2.5 Group-level ChannelGPT2 adapts generated data to individual subjects
@@
Up to this point, all trainings and analyses were done on MEG data from a single subject. We next looked at whether adding data from multiple subjects improves modelling and generation performance. This is in line with the overall goal of training such foundational forecasting models on multiple large datasets. Here we took a first step in exploring this by scaling ChanelGPT2 to the 15 subjects in the Cichy et al., (2016) data, which we refer to as ChanelGPT2-group. For adapting to multiple subjects and to capture variability over subjects, we used subject embeddings as described in Section 3. The main reason for only evaluating ChanelGPT2 on group data is the comparatively much poorer performance of Wavenet-based models in evoked timeseries generation.
@@
We were interested in whether the model generated evoked responses improved their similarity with the evoked responses from the real data, when using data from more subjects. To compare with the single-subject training we generated data using the subject embedding of that subject. The comparison of the evoked response of single-subject and group models for one 1 visual channel is shown in Figure 9. We found that generally ChannelGPT2-group produces evoked responses that are more smoothed than the single-subject model. This is possibly because the model learns to generate data that is closer to the average statistics over subjects, and while it can adapt its generation based on the subject label — this ability is not perfect.
Figure 9: Comparison of evoked responses in a visual channel (MEG2332) across single subject and group models. The stimulus onset is at 0 s and the stimulus offset is at 500 ms. Shading indicates 95% confidence interval across trials.
(a) Visual channel (MEG2332)
(b) Frontal channel (MEG0111)
@@
Figure 10: Comparison of evoked responses averaged across all subjects in the data (blue line) and the generated data from ChannelGPT2-group (orange line). The horizontal axis encompasses 1 second, stimulus onset is at 0 seconds and stimulus offset is at 0.5 seconds. Shading indicates 95% confidence interval across trials.
To test our hypothesis regarding ChannelGPT2-group generating more of an average across subjects, we generated data for all subjects (using appropriate subject embeddings) and compared the grand average evoked responses with those extracted from the MEG data of all subjects. Two channels are plotted in Figure 10. The evoked response averaged over all subjects is much noisier because of the high between-subject variability. However, we can see that indeed ChannelGPT2-group can generate this well, perhaps slightly smoother than the real data. Comparing these plots with Figure 9, it is also clear that it adapts its generation well to a specific subject compared to the group average.
A further way to test alignment between group-level evoked responses is to fit an HMM on the data of all subjects, and then infer state timecourses with this model on the generated data of all subjects from ChannelGPT2-group. By taking this approach we can directly match the evoked state timecourses between the real and generated timeseries. We trained an amplitude-envelope HMM (AE-HMM) with 6 states (Quinn et al.,, 2019) and show results in Figure 11. Two states that show strong activation during real task data show similar temporal signatures and amplitude changes in the generated data, albeit slightly noisier. In the generated data there are two additional states which seem to get activated during the trial. This indicates that while ChannelGPT2-group can capture some of the state-level dynamics, there is room for improvement.
(a) Data
(b) ChannelGPT2-group
Figure 11: Comparison of evoked state timecourses inferred from the data of all subjects and from the generated data of ChannelGPT2-group for all subjects. State indices are matched between the two plots, as the same fitted HMM model was used.
Finally, we examine the variability in state time courses over individual trials. For this we trained an 8-state HMM on the real data of a single subject, and inferred the state time-courses on both the single-subject ChannelGPT2 and ChannelGPT2-group generated data (with the appropriate subject embedding), obtaining matched states. We hypothesised that even if the average evoked responses are similar to the real data, GPT2 may not be able to generate trials with variability in the temporal activation of states. Figure 12 shows that this is indeed true for the single-subject ChannelGPT2 generated data. ChannelGPT2-group responses seem to include much higher temporal variability in state activations, though still falling short of the real data. This indicates that the model can capture some trial-to-trial variability through its exposure to multiple subjects, but has difficulty fully matching the complexity of real neural data. More data may be needed to improve this aspect of generation.
(a) Data
(b) ChannelGPT2
(c) ChannelGPT2-group
Figure 12: Comparison of trial-level variability in the evoked state timecourses of an HMM trained on real data and applied to the generated data of ChannelGPT2 and ChannelGPT2-group. Different colours represent different states (matched across models). Individual trials however are not matched and we cannot compare the plots at the trial-level, only as an aggregate visualisation of variability across trials.
2.6 Group-level ChannelGPT2 generates classifiable evoked responses
@@
We have shown that deep learning models, and particularly the channel-independent Transformer-based model (ChannelGPT2), can generate data with spatial, temporal, and spectral signatures similar to real data. We were next interested whether such a foundational model can aid in a downstream task. Specifically, we look at the ability of ChannelGPT2 to aid in the decoding of experimental task conditions in Cichy et al., (2016).
@@
We first investigated whether the task responses generated by the ChannelGPT2 model can be classified with performance comparable to trials of real data. This also further tests how well the model captures spatiotemporal task-related activity and information. The benefit of this approach is that if similar performance is obtained, then ChannelGPT2 could simulate an arbitrarily large number of trials to potentially improve decoding of real data through pretraining on this simulated data. This is a form of transfer learning where the decoding model, not the forecasting model, is transferred.
@@
We generated 20 trials for all 118 conditions for 1 subject with both ChannelGPT2 and ChannelGPT2-group. We trained separate linear neural network models on the real data (20 trials/condition) and the generated datasets, with an appropriate 4:1 train and validation set ratio. This achieved 17.6%, 1.9%, and 7.2% validation accuracy for the real data, ChannelGPT2, and ChannelGPT2-group, respectively. Thus the group model generates more classifiable subject-specific task-responses, but still does not reach the classifiability of real data. This and previous analyses indicate the group model successfully leverages larger datasets to produce more accurate task-related activity.
We also tried obtaining a decoding model directly from the ChannelGPT2-group forecasting model using Bayes’ theorem. We found limited 5% accuracy over 1 subject’s validation set (versus 40-50% with a discriminative decoder). This generative decoding approach may require larger datasets or more sophisticated architectures.
@@ 2.7 Transfer learning
A key advantage of generated data is the ability to generate huge amounts of surrogate data. We generated additional datasets with 40 and 60 trials/condition using ChannelGPT2-group. Training a decoder on these achieved 21.7% and 44.2% accuracy, respectively, exhibiting linear scaling of classification performance with simulated data amount. Critically, we assessed whether this simulated data can pretrain classifiers for transfer learning. We first pre-trained the neural network decoder on the 20-, 40-, and 60-trial generated datasets, then finetuned it (trained it further) on the real MEG dataset (20 trials/condition). As the simulated data used for pre-training increased, accuracy of the finetuned model improved rapidly. Zeroshot (no finetuning) performance on real data was above chance with 2%, 3%, and 4% accuracy, for increasing pretraining data quantities. Final accuracies after finetuning were 19.5%, 21.5%, and 23%, respectively. Thus, each additional 20 simulated trials/condition improved final decoding by ~2%. These results are summarised in Table 3.
Trained on (no. trials) Tested on MEG (20) Tested on GPT2 (same no. trial data)
MEG (20) 17.6 -
GPT2 (20) 2 7.2
GPT2 (40) 3 21.7
GPT2 (60) 4 44.2
GPT2 (20) + MEG (20) 19.5 -
GPT2 (40) + MEG (20) 21.5 -
GPT2 (60) + MEG (20) 23 -
Table 3: Summary of transfer learning results. The first column shows the data used for training the decoder, with the number of trials per condition shown inside the parenthesis. GPT2 refers to the ChannelGPT2-group generated data, while GPT (.) + MEG (20) is the fine-tuned decoder on the MEG data. The other two columns represent the validation data on which the decoder performance is shown. Accuracy values are provided in percentages. Chance level is $100/118$.
@@ 2.8 Ablation experiments
Ablation studies are a common approach in machine learning to understand model behaviour by selectively removing or altering components of the model (Meyes et al.,, 2019). We performed ablation experiments with ChannelGPT2 to investigate how well it can generate task-related brain activity under varied conditions without further training.
First, we evaluated the model’s ability to adapt to different trial durations. The results reported thus far are for a ChannelGPT2 trained on trials lasting 0.5 seconds. We generated data using the same fitted ChannelGPT2 model but with trial durations of 0.2 s and 0.8 s. As shown in Figure 13, ChannelGPT2 accurately adapted to the shorter and longer trials. The evoked responses matched the expected time-courses, with appropriate truncation or lack of second peaks due to stimulus offset. This demonstrates the model’s ability to generalise to varied trial durations despite being trained on a fixed duration.
@@
Next, we performed two experiments to determine whether ChannelGPT2 relies solely on timing information or also utilises the semantic content of the condition labels. First, we trained a model (ChannelGPT2-randomlabel) where the condition labels were shuffled randomly during training, breaking the semantic alignment between labels and evoked responses. Second, we trained a model (ChannelGPT2-1label) using a single condition label for all trials. This tests whether the model cheats by learning an average evoked response instead of adapting to each condition.
Figure 13: Evoked responses generated by ChannelGPT2 for trials of 0.2 s (orange), 0.5 s (blue), and 0.8 s (green). In all cases, the model was trained only on data containing trials of 0.5 s, but can then adapt appropriately to the different durations. The plotted channel is in the visual area (MEG2332).
(a) Channel 1 (MEG0111)
(b) Channel 2 (MEG1011)
(c) Channel 3 (MEG2332)
@@
Figure 14: Evoked responses for ChannelGPT2 models trained with shuffled or single condition labels, indicating reliance on semantic content. Three representative channels are presented. MEG0111 is anterior-left, MEG1011 is anterior-central, and MEG2332 is posterior-central. See main text for an explanation of model types. Stimulus onset is at 0 seconds, with stimulus offset at 0.5 seconds.
@@
As evident in Figure 14, both models failed to generate distinct evoked responses for different semantic conditions. This demonstrates that ChannelGPT2 leverages both timing and semantic information in the conditioning labels, rather than simply learning a stereotyped temporal template. Quantitatively, evoked response correlation with real data dropped to 44% and 56% for ChannelGPT2-randomlabel and ChannelGPT2-1label, respectively, compared to 74% for the full ChannelGPT2. Both the qualitative analysis and the correlation numbers indicate that ChannelGPT2-1label was somewhat closer to matching ChannelGPT2.
We also investigated the contributions of the channel and condition embeddings, by training two separate ablated models. As shown in Figure 15, removing the channel embeddings resulted in very similar PSD across channels in the generated data, indicating the model relies heavily on these embeddings to adapt generation per channel. The evoked responses in Figure 16 confirm that without channel embeddings, variability between channels is reduced. Removing the condition embeddings resulted in noisier power spectra of the generated data and no 20 Hz peak.
(a) Full ChannelGPT2
(b) No channel embedding
(c) No condition embedding
Figure 15: Generated power spectra for full ChannelGPT2 model (left) versus ablations. Both channel (middle) and condition embeddings (right) are critical for accurate spectral content.
(a) Channel 1 (MEG1011)
(b) Channel 2 (MEG1813)
(c) Channel 3 (MEG2332)
Figure 16: Comparison of generated evoked responses from ChannelGPT2 and the model with ablated channel embeddings (ChannelGPT2 no $W_{c}$) across 3 representative channels. Without channel embeddings the model fails to adapt evoked responses to different channels. The stimulus onset is at 0 seconds and the offset is at 0.5 seconds.
(a) t-SNE
(b) PCA
(c) Correlation
Figure 17: (a) (b) 2D projection of the channel embeddings from ChannelGPT2-group with t-SNE and PCA. Channels are coloured by their location on the scalp grouped into 5 major brain areas. (c) Plotting pairwise Euclidean distances of channels in real, physical space versus embedding space. Sensors that are near to each other in the real sensor montage tend to have more similar embeddings. Each point represents a different pair of channels. Correlation is 0.45.
Finally, we found that the channel embeddings encode spatial relationships, as sensors that are near to each other in the real sensor montage tend to have more similar embeddings. This is shown through a t-SNE and PCA projection of the embedding space in Figure 17. Correlation between pairwise Euclidean distances of channels in physical space and embedding space was 0.45 (Figure 17(c)).
3 Methods
3.1 Multi-channel Wavenet

### v2

@@ 5.3 Transformer generation better matches multivariate dynamics
Next, we assessed how well the fitted models can recursively generate data with the same spatial, temporal and spectral multi-channel characteristics as real MEG data. Hidden Markov Models (HMMs) are an established way for doing unsupervised discovery of multi-channel dynamics in real neuroimaging data, and have been used to characterise the spatial, temporal and spectral characteristics of brain networks in MEG data (Rabiner, 1989; Vidaurre et al., 2018a). All results in this section are obtained using a single sample subject.
@@
Figure 3: Distributions of dynamics summary statistics across the 12 states from an HMM inferred on real MEG multi-channel data from a single sample subject [left column], and from multi-channel data generated from 4 different forecasting models trained on the single sample subject. Four different summary statistics are shown describing the state dynamics (rows). ChannelGPT refers to GPT2MEG.
Separately, we trained each model on a single sample subject (with condition embeddings), generated multi-channel data, and then inferred a 12-state HMM, with each state modelling the multi-channel data as a multi-variate Normal (MVN) distribution. The number of HMM states was chosen based on previous work (Vidaurre et al., 2018a). Note that each recursively generated timeseries will be different, making detailed comparisons of the HMM state time courses across models meaningless. However, we can still consider broader differences in the generated dynamics, e.g. the speed of state switching. Note that since a separate HMM is trained each time, the states are not automatically matched between models or with the real data.
@@
We extracted four summary statistics on the different inferred state timecourses and compared their distributions over states. These are shown across models in Figure 3, alongside those for an HMM trained on the real multi-channel MEG data. Across the four summary statistics we can see that the real data has high variance in the distribution over states. AR(255) and WavenetFullChannelMix fail to produce data with variable state statistics, and even the mean over states is not captured well. WavenetFullChannel does a great job at capturing the mean of the state distributions, but still produces data with relatively invariant states. GPT2MEG is best at capturing the distributions across all four statistics, especially for the mean interval and switching rate. This shows that Transformer-based models can generate data that better matches the HMM-inferred dynamics of real MEG data.
5.4 Conditioned generation reproduces task-evoked responses
@@
Figure 4: Comparison of evoked timecourses of 2 channels across the task-conditioned models (via task-condition embedding); for real MEG data [blue line] from a single sample subject, and from data generated from the different forecasting models trained on the single sample subject. Stimulus onset is at 0 seconds and offset is at 500ms. The peak occurring after 500 ms indicates a visual response to the offset of the stimulus (removal of the image). Shading indicates the 95% confidence interval of the trial average. ChannelGPT refers to GPT2MEG.
The analyses in the previous section considered metrics for assessing the ability of the candidate autoregressive models to generate timeseries without requiring any a priori knowledge of the timing of brain activity. Here, we use knowledge of the experimental task timings in the Cichy et al. (2016) data, to provide insight into the ability of the models to generate realistic task data.
@@
By simple epoching of the generated timeseries based on the known task labels, we can compute evoked responses in the data generated by our models. We do this for all models except AR(255), as it is not able to straightforwardly include task labels in its model. To compare the shape of average evoked responses, we averaged over all epochs in both real data and the generated timeseries. This results in data of shape $\bar{\mathbf{X}}\in\mathbb{R}^{C\times T}$ where $C=306$ is the number of channels and $T=1000$ ms is the trial/epoch length.
@@
The evoked responses across our models and the real data in a frontal and a visual channel are shown in Figure 4. While Wavenet models completely fail to capture the evoked time-course, GPT2MEG does a remarkably good job, especially in the visual channel. This is not surprising as the dataset is collected from a visual experiment, so most activity is visual. GPT2MEG closely matches both the amplitude and the timing of the evoked response peaks across the whole 1-second epoch. Variability across trials is also well matched (as shown by the shading in Figure 4). These results generalize well to multiple subjects as will be shown in Section 5.5.
Full-sensor evoked-response correlation maps and HMM-state evoked analyses are provided in Appendix A.7.
5.5 Scaling to multiple subjects with subject embeddings
@@
Up to this point, all trainings and analyses were done on MEG data from a single sample subject. We next looked at whether combining data from multiple subjects improves modelling and generation performance. This is in line with the overall goal of training such foundational forecasting models on large datasets containing multiple subjects. Here we took a first step in exploring this by scaling GPT2MEG to the 15 subjects in the Cichy et al. (2016) data, which we refer to as GPT2MEG-group. For adapting to multiple subjects and to capture variability over subjects, we used subject embeddings (see Methods). The main reason for only evaluating GPT2MEG on group data is the comparatively much poorer performance of Wavenet-based models in evoked timeseries generation.
@@
We were interested in whether the model generated evoked responses improved their similarity with the evoked responses from the real data, when using data from more subjects. To compare with the single-subject training we generated data using the subject embedding of that subject. The comparison of the evoked response of single-subject and group models for one 1 visual channel is shown in Figure 5. We found that generally GPT2MEG-group produces evoked responses that are more smoothed than the single-subject model. This is possibly because the model learns to generate data that is closer to the average statistics over subjects, and while it can adapt its generation based on the subject label, this ability is not perfect.
@@
Figure 5: Comparison of evoked responses in a visual channel (MEG2332) for a single sample subject; using: real MEG data (blue), data generated from the GPT2MEG model trained on data from the single sample subject (orange), and data generated for the single sample subject (using an appropriate subject embedding) from the GPT2MEG-group model trained on all subjects (green). The stimulus onset is at 0 s and the stimulus offset is at 500 ms. Shading indicates 95% confidence of the trial mean. ChannelGPT refers to GPT2MEG.
Evoked responses across subjects and trial-level variability in evoked HMM state activations reported in A.11.
5.6 GPT2MEG-group generates classifiable evoked responses
@@
We have shown that the channel-independent, Transformer-based model (GPT2MEG), can generate data with spatial, temporal, and spectral signatures similar to real data. We next investigated whether we can use this as a foundational model in a downstream task. Specifically, we look at the ability of GPT2MEG to aid in the decoding of experimental task conditions in visual task dataset from (Cichy et al., 2016).
@@
We first investigated whether the task responses generated by the GPT2MEG model can be classified with performance comparable to trials of real data. This also further tests how well the model captures spatiotemporal task-related activity and information. Furthermore, if similar performance can be obtained, then GPT2MEG could be used to simulate an arbitrarily large number of trials to potentially improve decoding of real data through pretraining on the simulated data. This is a form of transfer learning, where the decoding model, not the forecasting model (e.g. GPT2MEG), is transferred.
@@
First, we generated 20 trials for all 118 conditions for one sample subject, using both GPT2MEG trained on the sample subject and GPT2MEG-group trained on all subjects (with the appropriate subject embedding of the chosen sample subject). We then trained separate linear neural network models on the real data (20 trials/condition) and these generated datasets, with an appropriate 4:1 train and validation set ratio. This achieved validation accuracies of 17.6% (real data), 1.9% (GPT2MEG), and 7.2% (GPT2MEG-group). In short, while the group model generates more classifiable subject-specific task-responses, it still does not reach the classification accuracy of real data. Nonetheless, this provides further evidence that GPT2MEG-group successfully leverages larger datasets to produce more accurate task-related activity.
@@ 5.7 Transfer learning
A key advantage of generated data is the ability to generate huge amounts of surrogate data. As in the previous section using 20 trials per condition, we generated additional datasets with 40 and 60 trials per condition using the GPT2MEG-group trained model. Training a decoder on these achieved validation accuracies of 7.2% (20 trials per condition), 21.7% (40 trials) and 44.2% (60 trials), exhibiting linear scaling of classification performance with the amount of simulated data.
Critically, we next assessed whether this simulated data can pretrain classifiers for transfer learning. First, we took the neural network decoder pre-trained on the 20-, 40-, and 60-trial generated datasets. We then finetuned the decoder (trained it further) on the real MEG dataset (20 trials per condition), and evaluated it on separate validation trials from the MEG data. As the quantity of generated data used for pretraining increased, accuracy of the finetuned model improved rapidly. Zeroshot (no finetuning) performance on real MEG data was above chance with 2% (20 trials per condition), 3% (40 trials), and 4% (60 trials) accuracy. Final accuracies after finetuning were 19.5% (20 trials), 21.5% (40 trials), and 23% (60 trials). Thus, each additional 20 GPT2MEG-group model-generated trials per condition improved final decoding by 2% on the real MEG trials. These results are summarised in Figure 6.
Figure 6: Summary of the decoding accuracies of visual stimuli when using different amounts of transfer learning. The horizontal axis represents which data the decoder was trained on. GPT2 (N) refers to the GPT2MEG-group generated data, while GPT2 (N) + MEG (20) is the fine-tuned decoder on the MEG data, where N is the number of trials per condition generated by GPT2MEG-group. The vertical axis shows the validation accuracy on the validation trials of the MEG data. Orange shows zeroshot performance, while with blue we denote the finetuned models. Chance level is $1/118$.
@@ 5.8 Ablation experiments
We performed ablation experiments with GPT2MEG for a single sample subject to investigate how well it can generate task-related brain activity under varied conditions without further training. In Appendix A.12 we show that it can accurately generate evoked activity for various stimulus lengths, while being trained only on the data with the original stimulus length.
@@
We performed two experiments to determine whether GPT2MEG relies solely on timing information or also utilises the semantic content of the condition labels. First, we trained a model (GPT2MEG-randomlabel) where the condition labels were shuffled randomly during training, breaking the semantic alignment between labels and evoked responses. Second, we trained a model (GPT2MEG-1label) using a single condition label for all trials. This tests whether the model cheats by learning an average evoked response instead of adapting to each task-condition.
(a) Ch. MEG0111
(b) Ch. MEG1011
(c) Ch. MEG2332
@@
Figure 7: Evoked responses for GPT2MEG models trained with shuffled or single condition labels, indicating reliance on semantic content. Three representative channels are presented for a single sample subject. MEG0111 is anterior-left, MEG1011 is anterior-central, and MEG2332 is posterior-central. See main text for an explanation of model types. Stimulus onset is at 0 seconds, with stimulus offset at 0.5 seconds. ChannelGPT refers to GPT2MEG.
@@
As evident in Figure 7, models with either with shuffled or single condition labels typically failed to generate distinct evoked responses for different semantic conditions. This demonstrates that GPT2MEG leverages both timing and semantic information in the conditioning labels, rather than simply learning a stereotyped temporal template. Quantitatively, evoked response correlation with real data dropped to 44% and 56% for GPT2MEG-randomlabel and GPT2MEG-1label, respectively, compared to 74% for the full GPT2MEG.
We further ablated the channel and condition embeddings and analyzed the learned channel-embedding geometry; these results indicate both embeddings are important and that spatial structure emerges in the channel embeddings (Appendix A.8).
6 Discussion
We introduced GPT2MEG, an autoregressive model for MEG trained by next-step prediction on unlabelled recordings. A simple discretization of MEG into tokens, together with channel, condition, and subject embeddings, allows a GPT-2–style Transformer (GPT2MEG) to generate realistic long-horizon dynamics and task-evoked responses. Across spectral, HMM-based dynamical, and evoked-response evaluations, Transformers outperform WaveNet variants and a linear AR baseline, despite similar one-step forecasting accuracy.
6.1 Limitations and future work
A core limitation of the channel-independent GPT2 model is that there is no direct leveraging of cross-channel information for each sensor prediction. We think that maintaining the innate inductive biases of Transformers, which emphasise 1D sequence modelling on embeddings of discrete tokens, is paramount. Different architectures, or more data, may enable proper utilisation of cross-channel dependencies. We tried various other approaches to mixing channel information beyond those reported, without success. For example, in the Wavenet model, we incorporated all channels in the input by concatenating embeddings, and for the GPT2 models, we tried mixing channels with convolutions. We tried concatenating the output of each channel and then predicting from this shared output using a different projection for each channel. We also attempted to increase receptive field, dropout, and model size.
Some of our findings substantiated that predicting the next timestep’s distribution (through the cross-entropy loss) may not serve as a robust measure of modelling performance. Future research should contemplate adopting multi-timestep or contrastive loss frameworks. A plausible strategy could involve deploying the VQ-VAE model across both channel and temporal dimensions, aiming to distill a coarser sequence of discrete tokens. Nevertheless, any quantisation-centric approach must carefully consider reconstruction error. We posit that a significant portion of the signal dynamics should be entrusted to the Transformer, given its adeptness in capturing complex dynamics.
A constraint in our modelling approach is its reliance on categorical task stimuli labels. Such an approach, while effective in our context, does not readily lend itself to scalability across diverse tasks and datasets. Robust representations tailored for various stimulus modalities, ranging from images to audio, can serve as conditioning embeddings. As shown by Défossez et al. (2022), tools such as wav2vec (Baevski et al., 2020) can be leveraged for this.
Future work should explore more flexible conditioning, study different self-supervised and transfer learning frameworks, and critically, apply similar analyses when scaling up across diverse, large electrophysiology datasets.
Acknowledgements
Impact Statement
This work develops autoregressive generative models for MEG that can simulate realistic task-evoked neural time series. Potential positive impacts include data augmentation for data-scarce decoding settings and reproducible benchmarking of analysis pipelines. Potential risks include misuse of synthetic data as a substitute for real measurements, amplification of biases present in the training data (e.g. demographic, acquisition, or preprocessing biases), and over-interpretation of generated signals as mechanistic evidence. We mitigate these risks by evaluating generation with multiple neuroscience-informed metrics. Future work should study calibration, dataset diversity, and safeguards for clinical and BCI applications.
Appendix A Appendix
A.1 Multi-channel Wavenet